# GMAI-VL & GMAI-VL-5.5M: A Large Vision-Language Model and A Comprehensive Multimodal Dataset Towards General Medical AI

### v1

@@ 9 Training Data Overview
Figure 6: Diagram of the three-stage training process.
10 Model Training Settings
Table 8: Training settings of GMAI-VL’s stage I, stage II, and stage III.
Settings Stage I Stage II Stage III
freeze LLM True True False
freeze MLP False False False
freeze Vision Encoder True False False
packing type soft packing soft packing soft packing
learning rate 1e-3 1e-4 1e-5
learning rate schedule cosine decay cosine decay cosine decay
optimizer AdamW AdamW AdamW
optimizer hyper-parameters $\beta_{1}=0.9,\beta_{2}=0.999$ $\beta_{1}=0.9,\beta_{2}=0.999$ $\beta_{1}=0.9,\beta_{2}=0.999$
input size 336x336 336x336 336x336
total batch size 32x8x2 32x4x4 32x4x4
drop rate 0.0 0.0 0.0
numerical precision DeepSpeed bf16 DeepSpeed bf16 DeepSpeed bf16
GPUs for training 32xA100 (80G) 32xA100 (80G) 32xA100 (80G)
The table. 8 presents the training settings for GMAI-VL across three stages, detailing key hyperparameters. The train diagram as illustrated in Fig. 6.
  • Stage I (Shallow alignment). During this stage, both the large language model (LLM) and the vision encoder are frozen, while the MLP remains trainable. The learning rate is set to $1e^{-3}$ with a cosine decay schedule, using AdamW as the optimizer. Input size is $336\times 336$, and the total batch size is $32\times 8\times 2$. Here, 32 refers to the number of GPUs used, 8 represents the micro-batch size per GPU, and 2 is the number of gradient accumulation steps. Notably, we use a soft packing technique, which allows each sample to actually contain multiple sequences—estimated at over 2 per sample.
  • Stage II (Deep alignment). The LLM remains frozen, while the vision encoder and MLP are unfrozen and trainable, allowing for deeper feature alignment. The learning rate is lowered to $1e^{-4}$, the total batch size is $32\times 4\times 4$, and other settings remain consistent with Stage I.
  • Stage III (Instruction tuning). In this stage, all components of the model are trainable, allowing for comprehensive fine-tuning. The learning rate is further reduced to $1e^{-5}$, the total batch size is $32\times 4\times 4$, while other parameters, including optimizer, remain unchanged across stages.
Each stage utilizes DeepSpeed for mixed-precision training (bf16) on 32 A100 (80GB) GPUs. To accelerate the training rate of the model, we utilize the soft packing method in xtuner***https://github.com/InternLM/xtuner, which packs multiple sequences into each sample to optimize memory usage and efficiency..
Figure 7: Examples of our model in various medical tasks, including image description, disease diagnosis, diagnostic report generation, free question answering, etc. The results demonstrate the model’s versatility across different imaging modalities and tasks, with support for both Chinese and English.
11 Results
Fig. 7 illustrates several examples of our model on various tasks including image description, disease diagnosis, diagnostic report generation, free question answering, etc. These examples cover a diverse range of medical imaging modalities and tasks, highlighting the model’s versatility and general capabilities within medical domain. Furthermore, GMAI-VL supports both Chinese and English, thereby facilitating bilingual functionality and expanding its applicability in diverse clinical settings.

### v2

@@ 6 Conclusion
In this part, we provide a comprehensive overview of all datasets used for training the GMAI-VL model. This includes the dataset names, categories, the amount of data used, and the proportion of training data allocated to each dataset during the three phases of model training. Fig. 4 visualizes the distribution of our training data, illustrating the proportion of each category and subcategory.
Table below summarizes the datasets employed, along with their respective categories and sizes. It is important to note that for some datasets, we performed data cleaning and bilingual translation, which may result in reported dataset sizes differing from the official numbers.
Part 3 - Results on the Validation and Test Sets of GMAI-MMBench for Clinical VQA Tasks
For further comparisons, Table 11 provides additional results, including a broader range of models, such as Open-Source LVLMs, Proprietary LVLMs, and Medical Special Models.
Table 11: Results on the val and test sets of GMAI-MMBench for clinical VQA tasks. The full names of the evaluated tasks can be found in Table 5 in literature[16]. The best model in each category is in red, while the second-best is blue.
Random Guess
Open-Source LVLMs
Proprietary LVLMs
Medical Special Model
Our Model
Part 4 - Quality Evaluation
Table 12: Prompt for Evaluating the Quality of Medical Image-Text Dialogues
You are a medical imaging expert tasked with scoring each medical text-image dialogue according to the following criteria. The score range is from 1 to 5, where a lower score indicates higher dialogue quality. Each dialogue should be evaluated based on the following criteria, considering language clarity, completeness, medical accuracy, and relevance to medical imaging. Below is an example of the dialogue content and related information:
- Modality: <image modality>(e.g., CT, MRI, X-ray, etc.)
- Department: <department>(e.g., Ophthalmology, Pulmonary etc.)
- Label: <label>(e.g., The pre-defined label belong to image)
- Dialogue Content:
 - Question: <question>Question content</question>
 - Answer: <answer>Answer content</answer>
Please score each dialogue based on the following dimensions:
1. Medical Information Accuracy:
 - Ensure that medical terminology, diagnostic processes, and treatment plans align with current medical standards and practices.
 - Assess whether there are any medical errors or misleading information that could impact diagnostic or treatment decisions.
 - Example: If an imaging description or medical terminology is incorrect, how does this affect the overall evaluation?
2. Language Clarity and Fluency:
 - Evaluate whether the dialogue is concise, clear, and easy to understand.
 - Assess whether the language is fluent and natural, and if there are any overly complex sentence structures that might hinder information transmission.
 - Example: Are there any unnecessary words or overly complex structures that slow down the understanding of the content?
3. Dialogue Completeness:
 - Evaluate whether the dialogue includes complete medical information, covering the patient’s history, imaging results, diagnostic information, and any necessary details.
 - Assess whether any essential information is missing, omitted, or incomplete.
 - Example: If imaging results are not mentioned or the information is incomplete, does this affect the overall evaluation?
4. Medical Imaging Relevance:
 - Evaluate whether the analysis of medical imaging information is thorough and accurate.
 - Ensure that the imaging data is appropriately interpreted, the descriptions are clear and accurate, and they align with the medical diagnosis.
 - Assess whether the imaging data effectively supports or supplements other medical information.
 - Example: If the description of the imaging data is inaccurate or incomplete, does it compromise the reliability of the diagnostic outcome?
Scoring Criteria:
- 1: The dialogue is concise, accurate, and clear, fully conforming to medical standards. The imaging interpretation is appropriate, the information is complete and accurate, and it fully supports the medical diagnosis.
- 2: The dialogue is generally accurate, the language is clear, and the content is complete. There may be slight imperfections in expression or minor inaccuracies in imaging interpretation, but overall quality is high.
- 3: The dialogue is accurate and complete, but there may be slight room for improvement in certain areas. For example, the imaging interpretation could be clearer, or language expression could be slightly more refined. Overall quality is good and suitable for training data.
- 4: The dialogue contains some notable medical errors or unclear language that affects accuracy. The imaging interpretation is problematic, or it does not integrate well with other clinical information, leading to incomplete content and a reduction in quality.
- 5: The dialogue contains errors or is unclear, potentially misleading medical decisions. There is a lack of effective imaging interpretation, the information is incomplete, and the language is unclear, which compromises the medical diagnosis. The overall quality is poor.
Output Format:
Please output the results in JSON format as follows:
{”score”: ”Quality score”, ”comments”: ”the comments for the quality criteria”,}
To ensure the quality of the generated image-text pairs, we propose a systematic evaluation process based on a 5-level scoring system. The definition of each scoring level is outlined in the prompt (Table 12). The primary objective of this process is to assess the quality of image-text pairs from a collection of 219 generated datasets and 42 publicly available open-source datasets. To ensure the representativeness and scientific rigor of the evaluation, we adopt a sampling survey methodology. Specifically, we randomly select a subset of image-text pairs from each dataset for quality assessment, from which we infer the overall quality of the entire dataset. This sampling survey strategy enables us to efficiently and scientifically evaluate the quality of the image-text pairs, maximizing the reliability and accuracy of the results while significantly reducing the manual review workload.
The evaluation process is conducted in multiple stages:
  1. 1.
    Initial Quality Assessment: From each dataset, 30 random image-text pairs are selected for evaluation. The first step involves an initial quality screening using a multimodal large language model (such as GPT-4o) to automatically assess the quality of the pairs. The model utilizes a predefined prompt, as illustrated in Table 12, to assign preliminary quality scores based on the defined criteria (accuracy of medical information, clarity and fluency of language, completeness of the dialogue, and relevance to medical imaging).
  2. 2.
    Expert Review: Following the model’s initial assessment, the 30 image-text pairs are then reviewed by five medical imaging experts. These experts validate the preliminary scores and provide their independent judgments on the quality of the pairs, ensuring that the evaluation aligns with current medical standards and practices.
  3. 3.
    Final Scoring: After expert review, the final score for each dataset is determined based on the consensus of the expert evaluations. If the average score of the 30 image-text pairs from a dataset is less than or equal to 3 (on a scale of 1 to 5), the dataset is classified as having satisfactory quality.
Conclusion: After rigorous evaluation, 95% (247/261) of the datasets met the predefined quality standards and were deemed acceptable. Our analysis showed that most datasets generated by GPT using the given labels exhibited high quality, with only a few requiring prompt optimization and additional information to meet the desired standards. For low-quality datasets, we addressed the issues by refining the prompt design, improving label specificity, and incorporating additional contextual information to ensure the generated quality met the requirements. In contrast, 14 open-release multimodal medical datasets, such as, PMC-OA[43], quilt-1m[33], often extracted from research papers or manually converted, exhibited various issues such as incomplete dialogues, misalignment between images and text, overly brief responses, inclusion of non-medical images, and poor image clarity. Due to these limitations, only a portion of the open-source datasets met the required quality standards.
This systematic approach ensures that only high-quality image-text pairs are included for further use in medical imaging research, with clear documentation of the evaluation process and criteria.
{CJK}
UTF8gbsn